OpenStudy (anonymous):

solve the equation in the interval [0,2pi) sinx-2sin^2x=0

OpenStudy (anonymous):

Factorize using a common factor of sin x then.....

OpenStudy (anonymous):

ok i got sinx(1-2sinx)=0

OpenStudy (anonymous):

Excellent, you now have two things that multiply to give a result of zero .........

OpenStudy (anonymous):

sinx=0 sinx= 1/2

OpenStudy (anonymous):

top result . Do you use "exact values"?

OpenStudy (anonymous):

yes and thats what i really have trouble with

OpenStudy (anonymous):

Do you use a little triangle to help you remember exact values?

OpenStudy (sirm3d):

\[\sin x-2\sin ^{2}x=0\] factor sinx \[\sin x \left( 1-2\sin x \right)=0\] equate each factor to zero (a) \[\sin x=0\] or (b)\[1-2\sin x=0\] the solution to (a) is \[x=0,\pi\] the solution to (b) is \[x=\pi/6,5\pi/6\]

OpenStudy (anonymous):

like with the unit circle?

OpenStudy (anonymous):



OpenStudy (anonymous):

like this (but neater)

OpenStudy (anonymous):

ok i understand how do you get the second angle the \[\frac{ 5\pi }{ 6}\]

OpenStudy (anonymous):

Thats where the unit circle comes in. Sin is positive in two quadrants. more pictures help?

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):